# Query Rewriting via LLMs

### v1

@@ 7.1.2 Query Readability
Figure 10: Query execution times on TPC-DS 100GB
7.1.3 Individual Queries
The above results were for entire benchmarks. We now drill down into the performance at the granularity of individual queries. Due to space limitations, we focus only on the TPC-DS benchmark here.
Figure 8 compares the performance of LITHE and SOTA on each of the 65 FPR queries in TPC-DS – note that the speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale. We first observe that, gratifyingly, rewrites are indeed capable of providing dramatic speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
As another example, Figure 10 shows a rewrite where only LITHE achieves 2X speedup. Here, the two subqueries in the original query (each with 3 joins) are replaced by a common set of joins, and the aggregations are conditionally computed using CASE statements.
Second, even among the 16 PR queries of SOTA, there are instances (e.g. Q6, Q47) where LITHE delivers substantially faster speedups than SOTA. Overall, across these 16 queries, LITHE attains a GM of 28.7 as compared to SOTA’s GM of 19.5.
Finally, note that LITHE matched or improved on SOTA for all the FPR queries.
7.1.4 Query Execution Times.
Thus far, we had considered optimizer estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 10 shows this metric (on a $\mathbf{log_{10}}$ scale) for the original query, LITHE, and SOTA on a representative set of queries executed on the TPC-DS 100GB database.
We observe that substantial benefits are achieved, even exceeding order-of-magnitude improvements in some cases. Of course, as expected, these speedups may not directly match the optimizer-estimates because of the inherent imprecision in optimizer models. For example, the 5 orders-of-magnitude speedup estimated for Q41 decreases to a still respectable 3 orders-of-magnitude. The opposite also occurs: the 2 orders-of-magnitude speedup estimated for Q45 now becomes a huge 4 orders-of-magnitude.
7.2 Ablation Analysis
7.2.1 Components of LITHE
A natural question at this stage is the role of the various techniques in LITHE towards achieving its large performance benefits. This analysis is captured in Table 8 which lists the PR and MPR for each technique when invoked in isolation. The table also shows the cumulative number of PR and MPR when the different techniques are combined in the order of listing.
Table 8: Ablation Analysis
Isolated Cumulative
# PR # MPR # PR # MPR
Basic Prompts (Section 3) 19 13 19 13
Rules R1 — R4 (Section 4.1) 17 12 23 15
Rules R5, R6 (Section 4.2) 44 34 60 48
MCTS (Section 5) 65 65 65 65
We observe in the table that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) are capable of achieving maximum speedups for around a quarter (15/65) of the FPR query set. When the statistics-infused Rules (R5, R6) are added, this fraction jumps to about three-quarters (48/65). And finally, when MCTS is included, we reach full coverage of the FPR.
Table 9: Rewrite Overheads of LITHE and SOTA on FPR queries
Benchmark Avg. Time (min) Avg. Tokens Avg. Cost USD
LITHE SOTA LITHE SOTA LITHE SOTA
TPC-DS 4.6 1.7 22390 20076 0.055 0.050
DSB 8.7 4.0 22203 15699 0.055 0.039
ARCHER 2.3 0.6 7284 5465 0.018 0.013
JOB 6.1 1.3 22958 13692 0.057 0.034
StackOverflow 7.1 3.2 22634 12759 0.056 0.032
7.2.2 Database-Sensitive Prompts.
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads. When this experiment was conducted, the SpeedupGM dropped precipitously to a paltry 2.3, a far cry from the 30.6 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
7.3 Rewrite Overheads
Having established the ability of LITHE to deliver performance-beneficial rewrites, we now turn our attention to the overheads incurred in this rewriting process. Here, in addition to rewrite time, the monetary outlay for LLM inferencing also has to be considered.
7.3.1 Transformation Time
The average processing time per FPR query is shown in Table 9 for LITHE and SOTA over the various benchmarks. The good news is that the rewrites are typically identified in a few minutes even for highly complex queries. We hasten to note that this investment is likely to be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. As a case in point, for TPC-DS Q30 on a 10GB database, the original version took 17 minutes to run to completion, whereas the LITHE rewrite took 7 minutes to transform but finished in less than a second! Further, when the database size was increased to 100GB, the LITHE rewrite executed in just 8 seconds, while the original version took more than an hour. Finally, many applications tend to use a set of canned queries which are run thousands of times. Thus, a relatively small one-time investment can be easily recovered over repeat executions of such queries.
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down into the various components of LITHE is shown in Table 10 for the TPC-DS queries. We observe here that the lion’s-share is taken by the initial prompt ensemble and the final MCTS module, largely arising from their “brute-force” combinatorial construction. But both these modules can potentially be improved – the choice of prompt via a classifier, and MCTS by early pruning of ill-formed (semantic/syntactic errors) queries. These options are discussed below in Section 7.3.3.
Table 10: LITHE Transformation Time Analysis
Time/Query (min) Tokens / Query
Avg. Max. Avg. Max.
Basic Prompts (Section 3) 1.2 5.2 2407 17778
Rules R1 — R4 (Section 4.1) 0.8 3.6 3076 11988
Rules R5, R6 (Section 4.2) 0.5 2.1 2114 5865
MCTS (Section 5) 2.1 10.6 14792 55055
7.3.2 Monetary Outlay
The average number of LLM tokens required by LITHE and SOTA for the various benchmarks, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are shown in Table 9. The good news is that the inference charges per query are just a few cents, making rewriting practical from a deployment perspective.
We also see that LITHE can be a factor of two more expensive than SOTA. To understand the source of this expense, the token costs incurred by each module of LITHE are shown in Table 10. Again, the lion’s share is taken by the MCTS module. This is to be expected due to its exploratory nature which can result in traversal of several paths in the tree – however, the additional expense appears to be justified by its output of significantly better rewrites. Finally, even this expense gets reduced when the time overheads are cut down, as described below.
7.3.3 Overheads Reduction
We design an LLM based classifier to accomplish this as follows. The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied. For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision. Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 40 percent, and the tokens by about 20 percent. However, there is a price to be paid – the PR is reduced to 59 and the SpeedupGM comes down to 21.7. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
Pruning in MCTS. A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output. In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early. We are currently working on the design and implementation of such checks.
7.4 Alternative Platforms (DB/Engine/LLM/T2S)
7.4.1 Anonymized Database
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of the GPT-4o LLM having already been trained well on our evaluated benchmarks, all of which are in the public domain. To investigate this issue, we created an anonymized version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their content. We then constructed rewrites for the 65 FPR queries (after syntactic changes to reflect the new schema) on this version. The results are shown in Table 12 for SOTA and LITHE. We observe that while the SpeedupGM marginally decreases, the performance profiles remain similar to those seen earlier for native TPC-DS, testifying to the robustness of the rewrite approaches.
Table 11: Impact of Classifier (TPC-DS)
Metrics No Classifier With Classifier
# PR 65 59
SpeedupGM 30.6 21.7
Avg. Tokens 22390 18296
Avg. Time 4.6 min 2.8 min
Table 12: Rewrite Performance on Anonymized Database
Approach # PR SpeedupGM
TPC-DS Masked TPC-DS Masked
SOTA 16 14 2.1 1.8
LITHE 65 64 30.6 29.4
7.4.2 Commercial DBMS
Another legitimate question could be whether the rewrites made amends for the relative lack of sophistication in the PostgreSQL optimizer, and that its benefits would disappear in premier database engines. To evaluate this issue, we performed TPC-DS rewrites on two popular and highly engineered commercial DBMS.
Table 13 compares the performance of LITHE and SOTA on these systems. Compared to PostgreSQL, the PR reduced considerably from 65 to 15 and 29, respectively, in case of LITHE, while SOTA had just 6 and 14 PRs respectively. Nevertheless, a significant fraction of the benchmark queries are still represented, providing a very respectable SpeedupGM of 5.3 and 3.4. So, LITHE has a role to play even on high-end database engines. From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
Table 13: Rewrite performance on Commercial Database Engines
Approach # PR SpeedupGM
Opt A Opt B Opt A Opt B
LITHE 15 29 5.3 3.4
SOTA 6 14 2.2 2.5
7.4.3 LITHE on LLaMA
In our concluding experiment, we evaluate the performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which is said to have several hundred billion parameters [5]. To attain practical inference times, the model was loaded using a low 4-bit quantization. Further, to ensure reproducibility and deterministic answers, the do_sample parameter was set to False, which forces the LLM to perform greedy decoding. In order to make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
For this environment, Table 14 shows the PR and SpeedupGM obtained on the micro-benchmark with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (see Tables 2 and  5), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that even smaller models can be fruitfully used in real-world environments.
Table 14: Micro-benchmark Performance with LLaMA
# PR SpeedupGM
LLaMA without MCTS 5 2.6
LLaMA with MCTS 9 7.2
7.4.4 Text-to-SQL
As mentioned in the Introduction, we could alternatively construct the SQL-to-SQL rewrite problem as a two-stage pipeline: SQL-to-Text followed by Text-to-SQL. To assess this option, we first implemented SQL-to-Text on GPT-4o with the prompt asking for conversion of the SQL query to a clear business description. Then, two open tools were evaluated for Text-to-SQL – CHESS [28] and SQLCoder [3], which have performed very well on the BIRD [17] and SQL-Eval [2] benchmarks, respectively. However, in all our experiments, the outcomes were uniformly poor with the rewrites often suffering semantic errors, and even among the few correct rewrites, there was not a single PR. Therefore, this alternative architecture does not appear to be a viable option.
@@ 8 Related Work
Query Booster [7] implements human centred rewriting – it provides an interface that the user can use to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query. There also exist other rule-based rewrite approaches that are designed for specific type of rewrites such as optimizing correlated window aggregations [32] and common expression elimination [9].
@@
A limitation, however, is that LLM-generated rewrite rules often fail to generalize beyond specific query pairs. Even when generalized rules are present, without accompanying examples, LLMs can struggle to apply rules correctly. Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
LLMs for Text-to-SQL. LLMs have also been extensively used for Text-to-SQL transformations [17, 33, 2, 3, 28]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [29, 23, 28]. On the other hand, the goal of SQL rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
@@ 9 Conclusions and Future Work
We investigated how the latent power of LLM technologies can be productively materialized in the context of SQL-to-SQL rewriting. Our study progressively infused database domain knowledge, such as redundancy removal rules and schematic+statistical metadata, into the LLM prompts. Further, the output telemetry of LLMs, in the form of token probabilities, was used to signal situations where the LLM lacked confidence, triggering exploration of a larger search space. Finally, a combination of logic-based and statistical tests was employed to verify the equivalence of the rewrites. An empirical evaluation over common database benchmarks showed that rewriting is a potent mechanism to improve query performance. The speedups obtained are significant and proof-of-concept, but there remains scope for improvement. For instance, while a speedup of 83X was produced for Q95 of TPC-DS, we were able to hand-craft an alternate rewrite with 99X speedup by moving a self-join from a CTE to the main query, thereby reducing scans of the web_sales table. We are currently working on adding a few select rules to implement such optimizations within the LITHE framework.
Our focus here was primarily on prompting-based strategies. In our future research, we also plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
@@
A1 : Detailed MCTS Algorithm
@@
A2 : Query Equivalence Testing
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g. Cosette[10], SQL-Solver [13], QED [30]). In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
To probabilistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples. The goal is to reduce the likelihood of non-equivalent queries returning the same results.
  1. 1.
  2. 2.
  3. 3.
3. Result Equivalence on the Entire Database. Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself. However, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
A3 : Examples used in Rules 1–6
@@
R6: Pre-filter fact tables in a CTE using those dimension tables having low selectivity values. Retain dimension table filters in main query. Do not create explicit join statements.
@@
Select item.id, item.code, item.price, country.name
from item, element, country, colour
where element.sourceid = item.sourceid
  and item.code = country.code
  and item.colorcode = colour.colorcode
  and element.zip > 8000
  and country.name = ’US’
  and colour.colour_name = ’Red’
order by item.id;
@@
( 1 ) element.zip > 8000 :: 0.002
( 2 ) country.name = ’US’ :: 0.4
( 3 ) colour.colour_name = ’Red’ :: 0.01
@@
with filtered_item as(
  Select *
  from item
  where
    item.sourceid in (
      Select
        element.sourceid
      where
        element.zip > 8000
    and item.colorcode in (
      Select
        colour.colorcode
      from
        colour
      where
        colour.colour_name = ’Red’
Select item.id, item.code, item.price, country.name
from filtered_item as item, element, country, colour
where
  item.sourceid = element.sourceid
  and item.code = country.code
  and item.colorcode = colour.colorcode
  and element.zip > 8000
  and country.name = ’US’
  and colour.colour_name = ’Red’
order by item.id;

### v2

@@ 8.2 Rewrite Migration to Optimizer
On the flip side, there appears to be an “impedance mismatch” against such integration for certain classes of rewrites. For example, consider the TPC-DS Q90 rewrite in Figure 9. The original query individually computed AM (morning) sales and PM (evening) sales, which were then used to compute the AM to PM ratio. The rewrite, however, extracted all relevant rows in one shot and computed the ratio using CASE statements – encoding such transformations as generic rules in the optimizer appears challenging, given the combinatorial ways in which such transformations can occur.
Therefore, a fruitful area of future research could be achieving a middle-ground between the disparate world-views of LLMs and traditional optimizers.
8.3 Revisiting Optimizer Plan Costing
As highlighted in Section 7, there were instances of substantive differences between the promised speedup and that delivered at run-time. In fact, to the extent that speedups could even turn out to be regressions! This is due to the brittleness of optimizer plan costing in the new spaces explored by the LLM. As a case in point, during our LITHE design process, the prefiltering in Rule R6 (Table 1) had initially not been restricted to self-joins. It resulted in the number of CPRs (on PostgreSQL) being as high as 65, with an astonishing CSGM of 30.6! However, upon execution, most of these rewrites turned out to be regressions, which led to our inclusion of the restriction. But note that we are artificially modulating the rewrite, rather than fixing the plan costing module, which is the principled solution. In sum, while plan cost modeling has been a long-standing area of research, there is now even more reason given the new rewrite spaces to study this topic further – for instance, the operator cost model could be extended using calibration techniques similar to those advocated in [40], while the operator cardinality model could be improved with attention-based techniques [23].
8.4 Agentic LLMs for Query Rewriting
An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment. This would allow a “LITHE agent” to leverage a memory store to log one or more of: (1) query execution times; (2) memory usages of slow queries; (3) prior interactions with users; and (4) useful/new rewrite rules learnt over time. Additionally, the agent could be given database engine access so that it can directly request metadata such as query execution plans, database statistics, live query analyzers, etc. By following a structured workflow – planning, acting, observing, and refining – the agent would then iteratively enhance query performance. This adaptive workflow supports continuous learning, so that query rewriting can become more efficient over time.
8.5 Scope of Semantic Equivalence Tools
As seen in the experiments section, the coverage of logic-based query equivalence testing on industrial-strength queries is rather limited. On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness. This limitation restricts the use of LITHE in a fully automated scenario, i.e., as a direct preprocessor to the query engine. Therefore, a key challenge is to improve logic-based coverage.
@@ 9 Related Work
Query Booster [8] implements human-centered rewriting – it provides an interface to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query. There are also rule-based rewrite approaches designed for specific types of rewrites such as optimizing correlated window aggregations [39] and common expression elimination [10].
@@
A limitation, however, is that LLM-generated rewrite rules often fail to generalize beyond specific query pairs. Even when generalized rules are present, LLMs can struggle to apply rules correctly if not provided with accompanying examples. Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
LLMs for Database Modules. LLM technologies have been advocated for a variety of database modules. For instance, they have been extensively used for Text-to-SQL transformations [22, 41, 2, 3, 33, 46]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [35, 29, 33]. On the other hand, the goal of S2S rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
In recent times, LLMs have also been considered for plan-hinting [7], join-order optimization [34], index selection [46], data pipelines [18], data management [21], etc. An LLM-enabled multi-modal query optimizer [37], where data spans text, image, audio and video domains, has also been proposed recently. These approaches can be used in conjunction with the LITHE system since they address orthogonal segments of the query processing pipeline.
@@ 10 Conclusions and Future Work
We investigated how the latent power of LLM technologies can be productively materialized in the context of SQL-to-SQL rewriting. Our study progressively infused database domain knowledge, such as redundancy removal rules and schematic+statistical metadata, into the LLM prompts. Further, the output telemetry of LLMs, in the form of token probabilities, was used to signal situations where the LLM lacked confidence, triggering exploration of a larger search space. Finally, a combination of logic-based and statistical tests was employed to verify the equivalence of the rewrites.
An empirical evaluation over common database benchmarks showed that rewriting is a potent mechanism to improve query performance. In fact, even order-of-magnitude speedups were routinely achieved with regard to both abstract costing and execution times. However, our results also showed a significant semantic distance between foundation models and query optimizers, with regard to both scope and precision, which would have to be bridged to fully leverage the latent power of LLMs. Further, our focus here was primarily on prompting-based strategies – in our future research, we plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
@@
Appendix A Rewrite Quality (Other Benchmarks)
We also evaluated LITHE on the following set of benchmarks.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Note that the queries in these workloads (with the exception of DSB) are mostly fast running (i.e. completing in less than 10 seconds). Therefore, with the primary goal to test the coverage of LITHE over a diverse variety of workloads (and not the deployment angle), we include all the queries in our analysis. The number of CPRs produced and the cost speedup delivered by LITHE and SOTA over these benchmarks are shown in Table 10.
Table 10: Comparing LITHE with SOTA on different benchmarks.
Benchmark # CPR CSGM
LITHE SOTA Union LITHE SOTA
DSB 9 3 9 7.7 1.7
ARCHER 22 19 22 2.1 1.95
JOB 4 2 4 1.9 1.4
StackOverflow 2 1 2 8.7 7.5
In case of DSB, LITHE produces CPR for 9 queries resulting in a highly productive CSGM of 7.7. Similar to the case with TPC-DS, LITHE performs better than SOTA both with respect to its coverage, as well as the cost speedups.
Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of CPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting. Nevertheless, LITHE continues to achieve better CPR coverage, whereas SOTA misses quite a few opportunities. Further, the CSGM of LITHE is visibly better than SOTA.
Appendix B Overheads Reduction for LITHE
Rule Selection Classifier
We design an LLM based classifier to accomplish this as follows: The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied. For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision. Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
Table 11: Impact of Classifier (TPC-DS)
Metrics Without Classifier With Classifier
# CPR 26 23
CSGM 11.5 8.5
Avg. Tokens 18427 16003
Avg. Time 5 min 2.4 min
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent. However, there is a price to be paid – the CPR is reduced to 24 and the CSGM comes down to 8.5. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
Pruning in MCTS
A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output. In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early. We are currently working on the design and implementation of such checks.
Appendix C Query Equivalence Testing via Sampling
If the original query is not within the scope of the logic-based QED tool, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
To statistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples. The goal is to reduce the likelihood of non-equivalent queries returning the same results.
  1. 1.
  2. 2.
  3. 3.
Appendix D Detailed MCTS Algorithm
@@
Appendix E Examples used in Prompts for Rules 1–6
@@
R6: Pre-filter tables that are involved in self-joins and have low selectivities on their filter and/or join predicates. Remove any redundant filters from the main query. Do not create explicit join statements.
@@
with total_price_cte as (
    select item.id, colour.colorcode, sum(item.price) total_price
    from item, color
    where item.colorcode = colour.colorcode
    group by item.id, colour.colorcode
select t_sec.id, t_first.colorcode
from total_price_cte t_first, total_price_cte t_sec
where t_sec.id = t_first.id
    and t_first.colorcode = ‘R’
    and t_sec.colorcode = ‘R’
    and t_first.total_price > 0
order by t_sec.id
limit 100;
@@
( 1 ) colour.colorcode = ‘R’ :: 0.01
@@
with total_price_cte as (
    select item.id, colour.colorcode, sum(item.price) total_price
    from item, color
    where item.colorcode = colour.colorcode
    group by item.id, colour.colorcode
filtered_total_price_cte as (
    select * from total_price_cte
    where colorcode = ‘R’
select t_sec.id, t_first.colorcode
from filtered_total_price_cte t_first, filtered_total_price_cte t_sec
where t_sec.id = t_first.id
    and t_first.total_price > 0
order by t_sec.id
limit 100;